Circular No. 6447

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

4U 1735–444 R. A. D. Wijnands and M. van der Klis, Astronomical Institute ‘An-

ton Pannekoek’, University of Amsterdam; J. van Paradijs, University of Amsterdam and University of Alabama in Huntsville; W. H. G. Lewin, Massachusetts Institute of Technology; F. K. Lamb, University of Illinois, Urbana/Champaign; B. Vaughan, California Institute of Technology; E. Kuulkers, Astrophysics Division, ESTEC, European Space Agency; and T. Augusteijn, European Southern Observatory, report: “RXTE observations of 4U 1735–444 performed on Aug. 1.60–1.68 UT show a single kHz quasiperiodic oscillation at frequency 1149 ± 3 Hz. The power spectral peak width is 30 ± 10 Hz and the root-mean-square amplitude 3.1 ± 0.3 percent (0.24–18.37 keV). This is the seventh example of kHz oscillations in a low-mass x-ray binary.”

CYGNUS X-1 S. N. Zhang, B. A. Harmon, and W. S. Paciesas report for the Compton

Gamma Ray Observatory BATSE team: “The blackhole candidate Cyg X-1 is experiencing a transition from a soft state back to its normal hard state. It went into the soft state in mid-May, following a sudden x-ray flux increase in the band 1.3–12 keV (IAUC 6404) and a rapid hard-x-ray-flux decrease in the band 20–200 keV (IAUC 6405). Since July 26, its 20- to 200-keV flux increased from ∼ 200 to ∼ 600 mCrab (as of Aug. 7). Its energy spectrum also evolved from a steep power law (photon index ∼ −2.5), without obvious break up to ∼ 200 keV, to a harder power law (index ∼ −2.0) with a break at ∼ 100 keV, quite similar to its hard-state spectrum (typically with index −1.8 and with a break at 100–150 keV). If its current trend of hard-x-ray-flux increase continues, Cyg X-1 will be back to its normal hard state in about a week.”

NOVA SAGITTARII 1996 Visual magnitude estimates, in part from the AAVSO: Aug. 5.86 UT,

10.8 (L. Monard, Pretoria, S. Africa); 6.88, 11.0 (J. Ripero Osorio, Madrid, Spain); 7.22, 10.8 (W. Dillon, Missouri City, TX); 7.93, 11.3 (P. Schmeer, Bischmisheim, Germany); 9.15, 11.2 (Dillon).

1996 August 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green